# A high resolution simulation of protoplanetary disk turbulence driven by the vertical shear instability

### v1

@@ Abstract
A high resolution fourth-order Padé scheme is used to simulate locally isothermal 3D disk turbulence driven by the vertical shear instability (VSI) using 268.4 M points. In the early non-linear period of axisymmetric VSI, angular momentum transport by vertical jets creates correlated N-shaped radial profiles of perturbation vertical and azimuthal velocity. This implies dominance of positive perturbation vertical vorticity layers and a recently discovered angular momentum staircase with respect to radius ($r$). These features are present in 3D in a weaker form. The 3D flow consists of vertically and azimuthally coherent turbulent shear layers with small vortices with all three components are active. Previously observed large persistent vortices in the interior of the domain driven by the Rossby wave instability are absent. We speculate that this is due to a weaker angular momentum staircase in 3D in the present simulations than in a previous simulation. The turbulent viscosity parameter $\alpha(r)$ increases linearly with $r$. At intermediate resolution, the value of $\alpha(r)$ at mid-radius is close to that of a previous simulation. The specific kinetic energy spectrum with respect to radial wavenumber has a power law region with exponent $-1.84$, close to the value $-2$ expected for shear layers. The spectrum with respect to azimuthal wavenumber has a $-5/3$ region and lacks a $-5$ region reported in an earlier study. Finally, it is found axisymmetric VSI has artifacts at late times, including a very strong angular momentum staircase, which in 3D is present weakly in the upper layers of the disk.
@@ 1.1 General introduction and motivation
A question posed by astronomical observations of protoplanetary disks (PPDs), the history of our own solar system, and the results of theoretical and numerical modeling is: what types of flow, temperatures, and magnetic field configurations occur at different locations and during different periods of disk evolution and how do they impact the dynamics and chemistry of solid particles which eventually form planets? The traditional motivation for studying PPD turbulence was to explain the so-called “anomalous viscosity” that gives rise to an accretion flow into the central star. More recently, however, the study of PPD turbulence is also very much motivated by the desire to explore its role in the pathway from micron sized grains to asteroid-sized bodies, called planetesimals, that are believed to be the progenitors of rocky planets, the putative solid cores of gas giant planets, and whose remnants are today’s asteroids, comets, and Kuiper belt objects. This motivation is reflected in a wide ranging 19-author review (Lesur et al., 2022) examining various aspects of disk turbulence both in isolation and in interaction with particles and grains embedded in the gaseous nebula.
@@ 1.2 Physics of VSI and previous theoretical work
The present work employs cylindrical coordinates $(z,r,\phi)$ which form a right-handed system with unit vectors satisfying $\mathbf{e}_{\phi}=\mathbf{e}_{z}\times\mathbf{e}_{r}$. The spherical radius is denoted $R=\left(z^{2}+r^{2}\right)^{1/2}$.
@@
Since the (stabilizing) Brunt-Väisälä frequency $N_{z}$ increases more rapidly with $|z|$ than the vertical shear $\partial\Omega/\partial z$, surface modes in the upper regions of the disk are the first to be damped as the cooling time increases. They obtain a key criterion for global instability:
@@
Table 1: Resolution comparison between representative previous 3D simulations (without radiative transport) and the present one. $H_{0}$: scale height at the mid-radius of the computational domain. $p$: radial density exponent. $q$: radial temperature exponent. $h\equiv H_{0}/R_{0}$, the disk aspect ratio. $R$ and $r$ denote the radius in spherical and cylindrical coordinates, respectively. $\Delta R$ and $\Delta r$ denote the width of the radial domain and include sponge regions. The resolution per scale height ($H_{0}$) is assessed at the mid-radius of the computational domain. †: The radial mesh for this simulation is logarithmic in $R$ and the value $17.1$ represents an average resolution per $H_{0}$.
@@ 1.3 Previous numerical work
Table 1 lists parameters in some previous 3D simulations without radiative transport and compares them with the present one. The metric for numerical resolution given in the table, introduced by Manger & Klahr (2018), is the number of grid points per scale height $H_{0}$ at the mid-radius of the domain.
@@
There is a series of 3D Pluto simulations with radiative transport and stellar irradiation (Stoll & Kley, 2014, 2016). The primary results are: (i) Persistence of VSI with radiative transport. (ii) A peak in the $T_{r\phi}$ Reynolds stress away from the midplane. Stoll et al. (2017) performed vertically isothermal simulations and studied the existence of a meridional mean flow due to Reynolds stress gradients which we confirm in §4.9. Table 1 shows that these simulations are marginally resolved in the azimuthal direction, having $\approx 5$ grid points per scale height along the circumference at mid-radius. Flock et al. (2020) also performed Pluto simulations with radiative transport and stellar irradiation. Parameters were chosen to represent a typical T Tauri system. They employed a large number of grid points, $1024\times 512\times 2044$ ($n_{R}\times n_{\theta}\times n_{\phi}$), which gives 70 grid cells per scale height. Their main findings were that: (i) VSI is able to vertically loft 1 mm grains such that their scale height ratio $H_{\mathrm{grains}}(r)/r=0.037$; this is much larger than the value $H/r=0.007$ for the HL Tau system, which is better fit by nonideal MHD simulations (Flock et al., 2017). (ii) A large persistent vortex is formed by the RWI at a radial location near the inner radial boundary, where the thermal relaxation times becomes low enough that VSI becomes viable. This causes a peak in the turbulence $\alpha(r)$ parameter, which in turn leads to a dip in surface density and a maximum in the RWI indicator function $\mathcal{L}(r)$.
@@
Here $q$ is the radial temperature exponent such that $T(r)\propto(r/r_{0})^{q}$, $H_{0}$ is the scale height at mid-radius ($r_{0}$), and $H(r)$ is the local scale height (see §2.1). To obtain the above stated metric (at mid-radius), multiply the tabulated number of grid points per scale height in the Table 1 by $0.31$ for the $h\equiv H_{0}/r_{0}=0.10$ cases, and by $\approx 0.10$ for the $h=0.03$ case. Doing this implies that for $h=0.10$, we have $22.9$ points in the meridional plane per $\lambda_{r,\mathrm{max}}(r_{0})$, while M20 have $5.7$ points. For their $h=0.03$ case, M20 have $1.9$ points.
@@ 2.3 Simulation parameters
Table 2 lists simulation parameters. The number of grid points is $512\times 512\times 1024$ ($n_{r}\times n_{z}\times n_{\phi}$). We choose $r_{0}$ to be the mid-radius of the computational domain. The density and temperature exponents set to $p=-3/2$ and $q=-1$, respectively. The disk aspect ratio is chosen to be $h\equiv H_{0}/r_{0}=0.1$. These three parameters were chosen to be the same as for case p1.5h0.1 in M20.
@@ 3.1 Perturbation velocity
Figure 3b shows the azimuthal velocity perturbation, $\delta u_{\phi}$. As mentioned earlier, the local linear stability analysis of L&P shows that $\delta u_{\phi}$ has the same magnitude as $\delta u_{z}$ when tilt is neglected. The same is approximately true here.
Figure 3c shows that the product $\delta u_{z}\delta u_{\phi}$ is mostly $\gtrless 0$ for $z\gtrless 0$. This important fact is explained as follows. The specific angular momentum $j_{z}\equiv u_{\phi}r$ follows fluid elements for axisymmetric flow [^1] For $p=\rho c_{\mathrm{i}}^{2}(r)$, the baroclinic term does not enter the $\omega_{z}$ transport equation (B13) for axisymmetric flow, and therefore the circulation $\Gamma_{z}=j_{z}$ is unaffected by baroclinic torque.:
@@
Figure 4a was constructed to provide an alternate way of illustrating the fact, which a contour plot will soon make clear, that the perturbation ($\delta\omega_{z}$) of vertical vorticity,
@@
is dominated by thin positive layers. (Equation (25) is valid for axisymmetric flow only.) Figure  4a shows that the correlation between $\delta u_{\phi}$ (solid) and $\delta u_{z}$ (dashed) is more remarkable than revealed in Figure 3. It also shows that both velocity components have N shaped profiles such that regions of positive slope are much steeper than regions of negative slope. We note that in the lower half of the disk (not plotted), $\delta u_{z}$ has mirrored N waves while $\delta u_{\phi}$ has normal N waves. The azimuthal vorticity is defined as
@@
The staircase structure of $j_{z}(r)$ is shown as the solid line in Figure 4b. In the 3D case (solid blue line), this pattern is less intense and coherent. Klahr et al. (2023) and MF24a,b were the first to point out the formation of a $j_{z}$ staircase for axisymmetric flow. The N waves in $u_{\phi}(r)$ simply reflect this staircase structure.
@@ 3.2 Perturbation vorticity
In the upper disk, $\omega_{\phi}<0$ (blue) shear layers are dominant. Such layers are adjacent to a thin sliver of opposite sign (red) vorticity, followed by a diffuse red region in between the shear layers. The opposite is true in the lower disk. These patterns can also be seen in Figure 9 (lower right panel) of MFa24. The dominant sign pattern of $\omega_{\phi}$ correlates with the N waves in $\delta u_{z}(r)$ discussed in §3.1. (Keep in mind that since the $\phi$ axis points into the page, a positive $\omega_{\phi}$ corresponds to clockwise rotation in the page.) The shear layers display vortices created by the KH instability.
@@
The radial perturbation vorticity (Figure 5d) is the weakest of the three components. Positive values are dominant in the upper disk, the opposite being true for $z<0$.
@@
for $c=4.35$. The choice of this distribution was inspired by Jiménez (1996) who obtains a similar result for the inner part of the velocity gradient PDF in 2D isotropic turbulence and for a collection of 2D Gaussian core vortices. For large vorticity values in 2D isotropic turbulence, the analysis of Falkovich & Lebedev (2011) predicts exponential tails, which would show up as linear behaviour on a semi-log plot. It is unclear whether this is true in the present case for $\delta\omega_{\phi}$. The red curve in Figure 6 confirms that $\delta\omega_{z}$ is skewed toward positive values consistent with Figure 5(b). To confirm the sign pattern for $\omega_{\phi}$, we plotted $p(\operatorname{sgn}(z)\delta\omega_{\phi})$ (not shown to avoid clutter) and found it to be skewed toward negative values consistent with the color contour plot (Figure 5(a)).
@@ 3.4 Increase of wavelength with time
which equals $0.24H_{0}$ at $r=7.5H_{0}$, the midpoint of the interval to be plotted below. On the other hand, if we inspect Figure 5(b) we find that the radial wavelength is more than twice larger being about $0.59H_{0}$ at the same radial location. To investigate this further, we ran the same case with a white noise initial perturbation instead of a superposition of waves with random phases as described in §2.3. This change was made to eliminate any artifact that might be present due to the presence of subharmonics of $\lambda_{r,\mathrm{max}}$ in the wavy initial condition. We find that wavelength increase is also observed with a white noise perturbation.
@@
Figure 9 depicts profiles of $u_{z}(r)$ at two instants (at $z/H_{0}=-2$) near the left end of the domain where the growth rate is the largest. An increase of wavelength with time is visually apparent. A value for the wavelength was obtained as the distance between downward zero crossings. At $t/T_{0}=10.2$ (blue curve), the average of the three wavelengths in the interval plotted is $\lambda_{r}=0.24$ which equals the most linearly amplified wavelength at the midpoint of the interval. On the other hand, at $t/T_{0}=35.6$, the average of the two wavelengths is larger, namely, $\lambda_{r}=0.37$. A temporal increase of wavelength can occur either due to nonlinearity or non-constant coefficients in the linear phase, i.e., spatial variation of the basic state gradient. Our guess is that the wavelength increase is due to nonlinearity and is related to a putative inverse cascade in axisymmetric VSI discussed further in §5. Radial wavelength increase is also observed in the radiative hydrodynamic axisymmetric simulations of MF24b (their Figure 11).
@@
since their $q=-1$ and $H_{0}/R_{0}=0.054$. Hence their wavelengths are between a factor of 1.5 to 2.6 larger than the linear instability one. Curiously, their Figure 19 does not show any tendency for the wavelength to increase with time. This is at odds with the axisymmetric radiative hydrodynamic simulations of M24b and should be investigated.
@@ 4.1 Perturbation vorticity in a meridional plane
We now turn to the 3D simulation. Figure 10 shows components of the perturbation vorticity ($\delta\vec{\omega}$) in a meridional ($rz$) plane at $t/T_{0}=444.64$ during the period when the strength of the Padé filter has been reduced to $\epsilon_{\mathrm{filter}}=0.015$. As in the axisymmetric case, the strongest component is $\delta\omega_{\phi}$ (panel b) and consists of at least 6 turbulent shear layers (TSLs) that extend across the vertical extent of the disk; they are numbered from right to left and alternate in the dominant sign of $\delta\omega_{\phi}$. The predilection for negative layers for $z>0$ and positive layers for $z<0$ observed in the axisymmetric case (Figure 5(a)) is still noticeable. A more accurate description is that half of each TSL with predominantly $\delta\omega_{\phi}\lessgtr 0$ is more organized for $z\gtrless 0$, undergoes KH instability, and becomes disorganized or filamentary for $z\lessgtr 0$. This was also the true in the axisymmetric case.
The vertical component ($\delta\omega_{z}$, panel (c)) is the next strongest; however, its values are about $1/3$ as large as in the axisymmetric case (compare with Figure 5(b)). Inspection of the vertical velocity (not shown) indicates that each TSL forms the boundary of a vertical jet. Panel (c) depicts $\delta\omega_{z}$ with numerical labels placed at the same locations as in panel (b). One observes that $\delta\omega_{z}$ is dominated by positive values and exists on either the upper or lower half of each TSL. This makes $\delta\omega_{z}$ regions in the upper and lower halves staggered relative to each other. The explanation for this was given in §3.1. To the extent that the $\delta\omega_{z}$ TSLs remain coherent in the azimuthal direction (which is the case in the upper layers of the disk as will be seen later), they correspond to a change in angular momentum as shown earlier (Figure 4, blue line). This is much weaker than in the axisymmetric case. We will see later that some of these layers undergo Kelvin-Helmholtz type instability as predicted in the L&P analysis.

### v2

@@ Abstract
A high resolution fourth-order Padé scheme is used to simulate locally isothermal 3D disk turbulence driven by the vertical shear instability (VSI) using 268.4 M points. In the early non-linear period of axisymmetric VSI, angular momentum transport by vertical jets creates correlated N-shaped radial profiles of perturbation vertical and azimuthal velocity. This implies dominance of positive perturbation vertical vorticity layers and a recently discovered angular momentum staircase with respect to radius ($r$). These features are present in 3D in a weaker form. The 3D flow consists of vertically and azimuthally coherent turbulent shear layers containing small vortices with all three vorticity components active. Previously observed large persistent vortices in the interior of the domain driven by the Rossby wave instability are absent. We speculate that this is due to a weaker angular momentum staircase in 3D in the present simulations compared to a previous simulation. The turbulent viscosity parameter $\alpha(r)$ increases linearly with $r$. At intermediate resolution, the value of $\alpha(r)$ at midradius is close to that of a previous simulation. The specific kinetic energy spectrum with respect to radial wavenumber has a power law region with exponent $-1.84$, close to the value $-2$ expected for shear layers. The spectrum with respect to azimuthal wavenumber has a $-5/3$ region and lacks a $-5$ region reported in an earlier study. Finally, it is found that axisymmetric VSI has artifacts at late times, including a very strong angular momentum staircase, which in 3D is present weakly in the disk’s upper layers.
@@ 1.1 General introduction and motivation
A question posed by astronomical observations of protoplanetary disks (PPDs), the history of our own solar system, and the results of theoretical and numerical modeling is: what types of flow, temperatures, and magnetic field configurations occur at different locations and during different periods of disk evolution, and how do they impact the dynamics and chemistry of solid particles which eventually form planets? The traditional motivation for studying flow turbulence in PPDs was to explain the so-called “anomalous viscosity” that gives rise to an accretion flow into the central star. More recently, however, the study of PPD turbulence is also very much motivated by the desire to explore its role in the pathway from micron sized grains to asteroid-sized bodies, called planetesimals, that are believed to be the progenitors of rocky planets, the putative solid cores of gas giant planets, and whose remnants are today’s asteroids, comets, and Kuiper belt objects. This motivation is reflected in a wide ranging 19-author review (Lesur et al., 2022) examining various aspects of disk turbulence both in isolation and in interaction with particles and grains embedded in the gaseous nebula.
@@ 1.2 Physics of VSI and previous theoretical work
The present work employs cylindrical coordinates $(z,r,\phi)$ which form a right-handed system with unit vectors satisfying $\mathbf{e}_{\phi}=\mathbf{e}_{z}\times\mathbf{e}_{r}$, and $\mathbf{e}_{\phi}$ pointing into the page. The spherical radius is denoted $R=\left(z^{2}+r^{2}\right)^{1/2}$.
@@
Since the (stabilizing) Brunt-Väisälä frequency $N_{z}$ increases more rapidly with $|z|$ than the vertical shear $\partial\Omega/\partial z$, surface modes in the upper regions of the disk are the first to be damped as $t_{\mathrm{rad}}$ increases. They obtain a key criterion for global instability:
@@
Table 1: Resolution comparison between representative previous 3D simulations (without radiative transport) and the present one. $H_{0}$: scale height at the midradius of the computational domain. $p$: radial density exponent. $q$: radial temperature exponent. $h\equiv H_{0}/R_{0}$, the disk aspect ratio. $R$ and $r$ denote the radius in spherical and cylindrical coordinates, respectively. $\Delta R$ and $\Delta r$ denote the width of the radial domain and include sponge regions. The resolution per scale height ($H_{0}$) is assessed at the midradius of the computational domain. †: The radial mesh for this simulation is logarithmic in $R$ and the value $17.1$ represents an average resolution per $H_{0}$.
@@ 1.3 Previous numerical work
Table 1 lists parameters in some previous 3D simulations without radiative transport and compares them with the present one. The metric for numerical resolution given in the table, introduced by Manger & Klahr (2018), is the number of grid points per scale height $H_{0}$ at the midradius of the domain.
@@
There is a series of 3D Pluto simulations with radiative transport and stellar irradiation (Stoll & Kley, 2014, 2016). The primary results are: (i) Persistence of VSI with radiative transport. (ii) A peak in the $T_{r\phi}$ Reynolds stress away from the midplane. Stoll et al. (2017) performed vertically isothermal simulations and studied the existence of a meridional mean flow due to Reynolds stress gradients which we confirm in §4.9. Table 1 shows that these simulations are marginally resolved in the azimuthal direction, having $\approx 5$ grid points per scale height along the circumference at midradius. Flock et al. (2020) also performed Pluto simulations with radiative transport and stellar irradiation. Parameters were chosen to represent a typical T Tauri system. They employed a large number of grid points, $1024\times 512\times 2044$ ($n_{R}\times n_{\theta}\times n_{\phi}$), which gives 70 grid cells per scale height. Their main findings were that: (i) VSI is able to vertically loft 1 mm grains such that their scale height ratio $H_{\mathrm{grains}}(r)/r=0.037$; this is much larger than the value $H/r=0.007$ for the HL Tau system, which is better fit by nonideal MHD simulations (Flock et al., 2017). (ii) A large persistent vortex is formed by the RWI at a radial location near the inner radial boundary, where the thermal relaxation times becomes low enough that VSI becomes viable. This causes a peak in the turbulence $\alpha(r)$ parameter, which in turn leads to a dip in surface density and a maximum in the RWI indicator function $\mathcal{L}(r)$.
@@
Here $q$ is the radial temperature exponent such that $T(r)\propto(r/r_{0})^{q}$, $H_{0}$ is the scale height at midradius ($r_{0}$), and $H(r)$ is the local scale height (see §2.1). To obtain the above stated metric (at midradius), multiply the tabulated number of grid points per scale height in the Table 1 by $0.31$ for the $h\equiv H_{0}/r_{0}=0.10$ cases, and by $\approx 0.10$ for the $h=0.03$ case. Doing this implies that for $h=0.10$, we have $22.9$ points in the meridional plane per $\lambda_{r,\mathrm{max}}(r_{0})$, while M20 have $5.7$ points. For their $h=0.03$ case, M20 have $1.9$ points.
@@ 2.3 Simulation parameters
Table 2 lists simulation parameters. The number of grid points is $512\times 512\times 1024$ ($n_{r}\times n_{z}\times n_{\phi}$). We choose $r_{0}$ to be the midradius of the computational domain. The density and temperature exponents are set to $p=-3/2$ and $q=-1$, respectively. The disk aspect ratio is chosen to be $h\equiv H_{0}/r_{0}=0.1$. These three parameters were chosen to be the same as for case p1.5h0.1 in M20.
@@ 3 Axisymmetric simulation: early non-linear stage
For all color contour plots in the paper, pure white pixels corresponds to values that exceed the range of the color map.
@@ 3.1 Perturbation velocity
Figure 3(b) shows the azimuthal velocity perturbation, $\delta u_{\phi}$. As mentioned earlier, the local linear stability analysis of L&P shows that $\delta u_{\phi}$ has the same magnitude as $\delta u_{z}$ when tilt is neglected. The same is approximately true here.
Figure 3(c) shows that the product $\delta u_{z}\delta u_{\phi}$ is mostly $\gtrless 0$ for $z\gtrless 0$. This important fact is explained as follows. The specific angular momentum $j_{z}\equiv u_{\phi}r$ follows fluid elements for axisymmetric flow [^1] For $p=\rho c_{\mathrm{i}}^{2}(r)$, the baroclinic term does not enter the $\omega_{z}$ transport equation (B13) for axisymmetric flow, and therefore the circulation $\Gamma_{z}=2\pi j_{z}$ is unaffected by baroclinic torque.:
@@
Figure 4(a) was constructed to provide an alternate way of illustrating the fact, which a contour plot will soon make clear, that the perturbation ($\delta\omega_{z}$) of vertical vorticity,
@@
is dominated by thin positive layers. (Equation (25) is valid for axisymmetric flow only.) Figure  4(a) shows that the correlation between $\delta u_{\phi}$ (solid) and $\delta u_{z}$ (dashed) is more remarkable than revealed in Figure 3. It also shows that both velocity components have N shaped profiles such that regions of positive slope are much steeper than regions of negative slope. We note that in the lower half of the disk (not plotted), $\delta u_{z}$ has mirrored N waves while $\delta u_{\phi}$ has normal N waves. The azimuthal vorticity is defined as
@@
The staircase structure of $j_{z}(r)$ is shown as the solid line in Figure 4(b). In the 3D case (solid blue line), this pattern is less intense and coherent. Klahr et al. (2023) and MF24a,b were the first to point out the formation of a $j_{z}(r)$ staircase for axisymmetric flow. The N waves in $u_{\phi}(r)$ simply reflect this staircase structure.
@@ 3.2 Perturbation vorticity
In the upper disk, $\omega_{\phi}<0$ (blue) shear layers are dominant. Such layers are adjacent to a thin sliver of opposite sign (red) vorticity, followed by a diffuse red region in between the shear layers. The opposite is true in the lower disk. These patterns can also be seen in Figure 9 (lower right panel) of MF24a. The dominant sign pattern of $\omega_{\phi}$ correlates with the N waves in $\delta u_{z}(r)$ discussed in §3.1. The shear layers display vortices created by the KH instability.
@@
The radial perturbation vorticity (Figure 5(d)) is the weakest of the three components. Positive values are dominant in the upper disk, the opposite being true for $z<0$.
@@
for $c=4.35$. The choice of this distribution was inspired by Jiménez (1996) who obtains a similar result for the inner part of the velocity gradient PDF in 2D isotropic turbulence and for a collection of 2D Gaussian core vortices. For large vorticity values in 2D isotropic turbulence, the analysis of Falkovich & Lebedev (2011) predicts exponential tails, which would show up as linear behavior on a semi-log plot. It is unclear whether this is true in the present case for $\delta\omega_{\phi}$. The red curve in Figure 6 confirms that $\delta\omega_{z}$ is skewed toward positive values consistent with Figure 5(b). To confirm the sign pattern for $\omega_{\phi}$, we plotted $p(\operatorname{sgn}(z)\delta\omega_{\phi})$ (not shown to avoid clutter) and found it to be skewed toward negative values consistent with the color contour plot (Figure 5(a)).
@@ 3.4 Increase of wavelength with time
which equals $0.24H_{0}$ at $r=7.5H_{0}$, the midpoint of the interval to be plotted below. On the other hand, if we inspect Figure 5(b) we find that the radial wavelength is more than twice larger, being about $0.59H_{0}$ at the same radial location. To investigate this further, we ran the same case with a white noise initial perturbation instead of a superposition of waves with random phases as described in §2.3. This change was made to eliminate any artifact that might be present due to the presence of subharmonics of $\lambda_{r,\mathrm{max}}$ in the wavy initial condition.
@@
We find that wavelength increase is also observed with a white noise perturbation. Figure 9 depicts profiles of $u_{z}(r)$ at two instants (for $z/H_{0}=-2$) near the left end of the domain where the growth rate is the largest. An increase of wavelength with time is visually apparent. A value for the wavelength was obtained as the distance between downward zero crossings. At $t/T_{0}=10.2$ (blue curve), the average of the three wavelengths in the interval plotted is $\lambda_{r}=0.24$ which equals the most linearly amplified wavelength at the midpoint of the interval. On the other hand, at $t/T_{0}=35.6$, the average of the two wavelengths is larger, namely, $\lambda_{r}=0.37$. A temporal increase of wavelength can occur either due to nonlinearity or non-constant coefficients in the linear phase, i.e., spatial variation of the basic state gradient. Our guess is that the wavelength increase is due to nonlinearity and is related to a putative inverse cascade in axisymmetric VSI discussed further in §5. Radial wavelength increase is also observed in the radiative hydrodynamic axisymmetric simulations of MF24b (their Figure 11).
@@
since their $q=-1$ and $H_{0}/R_{0}=0.054$. Hence their wavelengths are between a factor of 1.5 to 2.6 larger than for linear instability (for the locally isothermal equation of state). Curiously, their Figure 19 does not show any tendency for the wavelength to increase with time. This is at odds with the axisymmetric radiative hydrodynamic simulations of M24b and should be investigated.
@@ 4.1 Perturbation vorticity in a meridional plane
We now turn to the 3D simulation. Figure 10 shows components of the perturbation vorticity ($\delta\vec{\omega}$) in a meridional ($rz$) plane at $t/T_{0}=444.64$ during the period when the strength of the Padé filter has been reduced to $\epsilon_{\mathrm{filter}}=0.015$. As in the axisymmetric case, the strongest component is $\delta\omega_{\phi}$ (panel (b)) and consists of at least 6 turbulent shear layers (TSLs) that extend across the vertical extent of the disk; they are numbered from right to left and alternate in the dominant sign of $\delta\omega_{\phi}$. The predilection for negative layers for $z>0$ and positive layers for $z<0$ observed in the axisymmetric case (Figure 5(a)) is still noticeable. A more accurate description is that half of each TSL with predominantly $\delta\omega_{\phi}\lessgtr 0$ is more organized for $z\gtrless 0$, undergoes KH instability, and becomes disorganized or filamentary for $z\lessgtr 0$. This was also true in the axisymmetric case.
The vertical component ($\delta\omega_{z}$, Figure 10(c)) is the next strongest; however, its values are about $1/3$ as large as in the axisymmetric case (compare with Figure 5(b)). Inspection of the vertical velocity (not shown) indicates that each TSL forms the boundary of a vertical jet. The numerical labels in panel (c) are placed at the same locations as in panel (b). One observes that $\delta\omega_{z}$ is dominated by positive values and exists on either the upper or lower half of each TSL. This makes $\delta\omega_{z}$ regions in the upper and lower halves staggered relative to each other. The explanation for this was given in §3.1. To the extent that the $\delta\omega_{z}$ TSLs remain coherent in the azimuthal direction (which is the case in the upper layers of the disk as will be seen later), they correspond to a change in angular momentum as shown earlier (Figure 4, blue line). This is much weaker than in the axisymmetric case. We will see later that some of these layers undergo Kelvin-Helmholtz type instability as predicted in the L&P analysis.